# Lower Bounds for Differential Privacy Under Continual Observation and Online Threshold Queries

### v1

@@ 1 Introduction
Typically one aims for a stronger privacy definition in which the neighboring datasets are not fixed in advanced, and can be adaptively chosen as the execution progresses. Since in this work we focus on negative results, then working with the above “non-adaptive” privacy definition only makes our results stronger.
@@ 1.2.2 Application II: Private online learning
Let $T\geq 100$ denote the number of time steps, and let $\mathcal{X}=\{0,1,2,\dots,T\}$ be the data domain. Let $\mathcal{H}$ be the class of point functions over $\mathcal{X}$. Let $\delta=\frac{1}{20\log(T)}$. Then, any $(\varepsilon,\delta)$-DP online learner for $\mathcal{X}$ that answers $T$ queries must make $\Omega(\log(T))$ mistakes with constant probability.
@@
Let $\mathcal{X}$ be a domain. The class of point functions over $\mathcal{X}$ admits an $(\varepsilon,\delta)$-DP online predictor with mistake bound $O(\frac{1}{\varepsilon}\log\frac{1}{\delta})$.
The proofs of Theorem 1.11 and Lemma 1.12 are given in the appendix.
@@ 2.2 The Formal Details
[…]
@@
1 Initialize $\ell\leftarrow 1$, $r\leftarrow T$, $D^{(0)}\leftarrow(0,0,\dots,0)$, and $c_{\mathrm{total}}\leftarrow 0$ for $i=1,\dots,k$ do
       $m\leftarrow\frac{1}{2}(\ell+r+1)$   // midpoint $m$ is always an integer
[…]
3             Set the $\ell$-th entry of $D^{(i-1)}$ to $1$ to obtain $D^{(i)}$ $c_{\mathrm{total}}\leftarrow c_{\mathrm{total}}+p_{\ell}^{D^{(i)}}$ $\ell\leftarrow\ell+1$ $r\leftarrow m-1$
@@
6             Set the $m$-th entry of $D^{(i-1)}$ to $1$ to obtain $D^{(i)}$ $c_{\mathrm{total}}\leftarrow c_{\mathrm{total}}+c_{\mathrm{left}}+p^{D^{(i)}}_{p}$ $\ell\leftarrow m+1$
@@ 3.2.2 The Key Lemma
  • First, We identify a subset $P^{(3)}$ of $P^{(2)}$ such that $\Pr[\mathcal{A}(S\cup\{x_{1},x_{2}\})\text{ outputs $\top$ for the $2$nd query}]$ is almost the same (i.e., up to an error of $\eta$) for every $S\subseteq P^{(2)}$.
@@ Appendix A Lower Bound for Private Online Learning
of Theorem 1.11   Suppose for the sake of contradiction that such a learner exists and denote it by $\mathcal{A}$. Denote $k=\sqrt{T}$, and consider the following distribution over query sequences.
  • For the $i$-th query when $i\in[1,k]$, send $(x_{i},0)$ (where $\textsf{Uniform}[a,b]$ is the uniform distribution on $\{a,\ldots,b\}$).
@@
that is, the average probability, over steps $i$ in the phase, of the probability that the released hypothesis is $1$ on $k^{\prime}$ is at least $0.001/k$.
@@
Appendix B JDP Mirror
Input: Parameters $\varepsilon<0.6,\delta$, $K\geq 1$
@@ Appendix B JDP Mirror
$K_{1}\sim P_{\delta}$; $K_{2}\sim P_{\delta}$   // independently, where $P_{\delta}$ is as in Claim 3
$C_{1},C_{2}\leftarrow 0$  // counters towards phase transition
for $t\in[T]$ do $r_{t}\sim\textsf{Bern}[\varepsilon]$  // independent step activation bits // Phase I:
for $t=1,\ldots$ do // Phase I
       Receive an input bit $\Delta_{t}\in\{0,1\}$ Output $y_{t}\equiv\bot$ if $\Delta_{t}=1\And r_{t}=1$ then // increment counter
            $C_{1}\leftarrow C_{1}+1$ if $C_{1}=K_{1}$ then  $t_{1}\leftarrow t+1$; break   // test and go to Phase II
@@
// Phase II:
for $t=t_{1},\ldots$ do // Phase II
       Receive an input bit $\Delta_{t}\in\{0,1\}$ if $\Delta_{t}\And r_{t}=1$ then
             Output $y_{t}=\top$ $C_{2}\leftarrow C_{2}+1$ if $C_{2}=K_{2}$ then  $t_{2}\leftarrow t+1$; break   // test and go to Phase III
      else
            Output $y_{t}=\bot$
       end if
end for
// Phase III:
for $t=t_{2},\ldots$ do // Phase III
       Receive an input bit $\Delta_{t}\in\{0,1\}$ if $\Delta_{t}=1$ then Output $y_{t}=\top$ else Output $y_{t}=\bot$
end for
@@
In this section we present a construction for (a variant of) the threshold monitor problem that satisfies a weaker variant of differential privacy known as joint differential privacy (JDP) and has an error that is independent of $T$. In the following section we will leverage this construction in order to present our private online predictor for point functions. We first recall the definition of JDP.
Definition B.1 (Kearns et al. (2015)).
@@
Let $\mathcal{M}:X^{n}\rightarrow Y^{n}$ be a randomized algorithm that takes a dataset $S\in X^{n}$ and outputs a vector $\vec{y}\in Y^{n}$. Algorithm $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-joint differential privacy (JDP) if for every $i\in[n]$ and every two datasets $S,S^{\prime}\in X^{n}$ differing only on their $i$th point it holds that $\mathcal{M}(S)_{-i}\approx_{(\varepsilon,\delta)}\mathcal{M}(S^{\prime})_{-i}$. Here $\mathcal{M}(S)_{-i}$ denotes the (random) vector of length $n-1$ obtained by running $(y_{1},\dots,y_{n})\leftarrow\mathcal{M}(S)$ and returning $(y_{1},\dots,y_{i-1},y_{i+1},\dots,y_{n})$.
We present an algorithm that satisfies JDP for the following Mirror problem. For a sequence of input bits $(\Delta_{t})_{t\in[T]}$ where $\Delta_{t}\in\{0,1\}$, the goal is to mirror the input, returning $y_{t}=\bot$ for $\Delta_{t}=0$ and returning $y_{t}=\top$ for $\Delta_{t}=1$. The constraint is that a $\top$ may be returned only after at least one prior $1$. We permit one-sided mistakes, that is, reporting $y_{t}=\bot$ when $\Delta_{t}=1$. The goal is to minimize the number of mistakes.
Algorithm 2 initializes by sampling independent $K_{1}$ and $K_{2}$ from the distribution $P_{\delta}$ as in Claim 3. For input $\Delta_{t}=0$, it always returns $y_{t}=\bot$. For input $\Delta_{t}=1$ the response depends on the phase. The algorithm has three internal phases. During the first phase, it returns $y_{t}=\bot$ for all steps. It maintains a counter $C_{1}$ that is incremented by an independent $r_{t}\sim\textsf{Bern}[\varepsilon]$ after each input with $\Delta_{t}=1$. When $C_{1}=K_{1}$, the algorithm internally transitions to the second phase. During the second phase, when $\Delta_{t}=1$ it again samples $r_{t}\sim\textsf{Bern}[\varepsilon]$ and increments a counter $C_{2}$ by $r_{2}$. It also returns $y_{t}=\top$ when $r_{t}=1$. When $C_{2}=K_{2}$, the algorithm internally transitions to the third phase. In the third phase, it always returns $y_{t}=\top$ when $\Delta_{t}=1$.
Claim 3 ($P_{\delta}$ distribution).
For any $\delta>0$, there is a discrete distribution $P_{\delta}$ on $\mathbb{N}$ with support $\{1,\ldots,L=O(\log(1/\delta)\}$ and such that $P(1)=P(L)=\delta$, $P(i)\approx_{0.5}P(i+i)$ for $i\in[L-1]$, mean $\mu=(L+1)/2$, and for $c\geq 0$, $\Pr[|X-\mu|>c=e-{\Theta(c)}$.
We show the following:
Lemma B.2 (Mistake bound of JDP-Mirror).
The expected number of mistakes made by Algorithm 2 is $O(\varepsilon^{-1}\log(1/\delta))$. The probability that the number of mistakes exceed $c\cdot\varepsilon^{-1}\log(1/\delta)$ is $\exp(-c\cdot O(\log(1/\delta))).$
Proof B.3.
During the first two phases, the number of $\Delta_{t}=1$ steps between two that result in incrementing the counter is $\textsf{Geom}[\varepsilon]$. These $\textsf{Geom}[\varepsilon]$ random variables are independent. During the first phase we incur mistakes also on the $K_{1}$ steps that are counted. During the second phase we do not incur mistakes on the $K_{2}$ steps that are counted. The distribution of the number of mistakes conditioned on $K_{1}$ and $K_{2}$ is therefore $\textsf{NegBinomial}(K_{1}+K_{2},\varepsilon)-K_{2}$. [^6] Recall that $\textsf{NegBinomial}(K,\varepsilon)$ is a sum of $K$ $\textsf{Geom}[\varepsilon]$ random variables. The number of mistakes is stochastically smaller than $\textsf{NegBinomial}(2L=O(\log(1/\delta),\varepsilon)$ since $K_{1}+K_{2}=O(\log(1/\delta)$ (by Claim 3). The expected value is therefore $O(\log(1/\delta/\varepsilon)$ and the concentration claim follows from a bound on the upper tail of the negative binomial distribution (4).
Theorem B.4 (JDP of JDP-Mirror).
Algorithm 2 is $(O(\varepsilon),O(\delta))$-JDP. That is, for any two adjacent sequences that differ in position $i^{*}$
$\boldsymbol{\Delta}$ $=\Delta_{1},\ldots,\Delta_{i^{*}-1},\Delta_{i^{*}}=1,\Delta_{i^{*}+1},\ldots,\Delta_{T}$
$\boldsymbol{\Delta}^{\prime}$ $=\Delta_{1},\ldots,\Delta_{i^{*}-1},\Delta_{i^{*}}=0,\Delta_{i^{*}+1},\ldots,\Delta_{T}$
and a set $Y$ of potential outputs
$\Pr[\texttt{JDPM}(\vec{\Delta})_{-i^{*}}\in Y_{-i^{*}}]\approx_{O(\varepsilon),O(\delta)}\Pr[\texttt{JDPM}(\vec{\Delta}^{\prime})_{-i^{*}}\in Y_{-i^{*}}].$
Remark B.5 (JDP-Delayed-Mirror).
JDP-Mirror (Algorithm 2) as described returns its first $y_{t}=\top$ after processing $\textsf{Geom}[\varepsilon]$ tops. Delayed-Mirror is an extension of JDP-Mirror where for an input parameter $K\geq 1$, the value $K-1$ is added to the sampled $K_{1}$ at initialization. The following holds with this modification:
  • The first $y_{t}=\top$ is returned only after $\textsf{NegBinomial}(K,\varepsilon)$ inputs with $\Delta_{t}=1$ are processed. From (5) and (4), for any $c>1$, the probability of the first $\top$” after fewer than $K/(c\cdot\varepsilon)$ or more than $cK/\varepsilon$ such inputs is $e^{-O(cK)}$.
@@
The privacy analysis (Section B.2) goes through with small modifications in the case $i^{*}<w_{1}$.
B.1 Notation and tools
Overloading the notation, for two nonnegative reals $a,b$ we write that $a\approx_{\varepsilon,\delta}b$ if $a\leq e^{\varepsilon}b+\delta$ and $b\leq e^{\varepsilon}a+\delta$. For two random variables $X,Y$ over domain $\mathcal{U}$, we write $X\approx_{\varepsilon,\delta}Y$ if for any $U\subset\mathcal{U}$, $\Pr[X\in U]\approx_{\varepsilon,\delta}\Pr[Y\in U]$. We will interchangeably use notation for random variables and distributions.
We will use the following tail bound for the Negative Binomial distribution Janson (2018): For $c>1$,
$\Pr_{X\sim\textsf{NegBinomial}(n,p)}[X>c\cdot n/p]$ $\leq e^{-O(cn)}$ (4)
$\Pr_{X\sim\textsf{NegBinomial}(n,p)}[X<\frac{1}{c}\cdot n/p]$ $\leq e^{-O(cn)}\ .$ (5)
Privacy Analysis via Excluded Runs
@@ B.1 Notation and tools
For two datasets $D$ and $D^{\prime}$ and $y$ we denote
$\Pr[\mathcal{M}(D)=y]\approx_{\varepsilon\mid F}[\mathcal{M}(D^{\prime})=y]$
if
$(1-F(\mathcal{M},D,y))\Pr[\mathcal{M}(D)=y\mid\neg F]\leq e^{\varepsilon}\Pr[\mathcal{M}(D^{\prime})=y]$
$(1-F(\mathcal{M},D^{\prime},y))\Pr[\mathcal{M}(D^{\prime})=y\mid\neg F]\leq e^{\varepsilon}\Pr[\mathcal{M}(D)=y]$
We use the notation $\mathcal{M}(D)\approx_{\varepsilon\mid F}\mathcal{M}(D^{\prime})$ if for all output $y$ it holds that $\Pr[\mathcal{M}(D)=y]\approx_{\varepsilon\mid F}[\mathcal{M}(D^{\prime})=y]$.
Lemma B.6 (Excluded Runs).
Let $\mathcal{M}$ be a randomized algorithm and $F$ be a predicate as above. Let $D$ and $D^{\prime}$ be such that
  • $F(\mathcal{M},D)\leq\delta$ and $F(\mathcal{M},D^{\prime})\leq\delta$
  • $\mathcal{M}(D)\approx_{\varepsilon\mid F}\mathcal{M}(D^{\prime})$
Then $\mathcal{M}(D)\approx_{\varepsilon,\delta}\mathcal{M}(D^{\prime})$.
Proof B.7.
Let $Y$ be a subset of the output domain
$\Pr[\mathcal{M}(D)\in Y]$ $=\sum_{y\in Y}\Pr[\mathcal{M}(D)=y]$
$=\sum_{y\in Y}(1-F(\mathcal{M},D,y))\cdot\Pr[\mathcal{M}(D)=y\mid\neg F]+F(\mathcal{M},D,y)\cdot\Pr[\mathcal{M}(D)=y\mid F]$
$\leq e^{\varepsilon}\sum_{y\in Y}\Pr[\mathcal{M}(D^{\prime})=y]+\sum_{y\in Y}F(\mathcal{M},D,y)$
$\leq e^{\varepsilon}\Pr[\mathcal{M}(D^{\prime})\in Y]+\delta\ .$
B.2 JDP of Algorithm 2
Fix two adjacent sequences:
$\boldsymbol{\Delta}$ $=\Delta_{1},\ldots,\Delta_{i^{*}-1},1,\Delta_{i^{*}+1},\ldots,\Delta_{T}$
$\boldsymbol{\Delta}^{\prime}$ $=\Delta_{1},\ldots,\Delta_{i^{*}-1},0,\Delta_{i^{*}+1},\ldots,\Delta_{T}$
let $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ respectively be the random variables that are the output of Algorithm 2 on $\boldsymbol{\Delta}$ and $\boldsymbol{\Delta}^{\prime}$. We need to show that
$\boldsymbol{y}_{-i^{*}}\approx_{O(\varepsilon),O(\delta)}\boldsymbol{y}^{\prime}_{-i^{*}}\ .$ (6)
Observe that positions $t\not=i^{*}$ with $\Delta_{t}=0$ must have $a_{t}=\bot$ and moreover, they have no effect on all other outputs algorithm. It therefore suffices to establish (6) on compressed sequences where these positions are removed. Going forward, we assume that $\Delta_{t}=\Delta^{\prime}_{t}=1$ for $t\not=i^{*}$ and $\Delta_{i^{*}}=1$ and $\Delta^{\prime}_{i^{*}}=0$.
We assume without loss of generality that the input includes at least $2L$ steps $t$ where $\Delta_{t}=1$ (and therefore Phase III is reached). When this is not the case we can pad the input and work with a larger $T$. This assumption simplifies the analysis.
We use the following:
Claim 4 (Geometric sequence).
For any $\varepsilon<0.6$, $i\geq 1$ and $X\sim P_{\delta}$ it holds that $\sum_{j\geq 0}\Pr[X=i+j]\varepsilon^{j}=O(\Pr[X=i])$

### v2

@@ 1 Introduction
Typically one aims for a stronger privacy definition in which the neighboring datasets are not fixed in advanced, and can be adaptively chosen as the execution progresses. Since in this work we show negative results, then focusing on the above “non-adaptive” privacy definition only makes our results stronger.
@@ 1.2.2 Application II: Private online learning
Let $T\geq 100$ denote the number of time steps, and let $\mathcal{X}=\{0,1,2,\dots,T\}$ be the data domain. Let $\mathcal{H}$ be the class of point functions over $\mathcal{X}$. Let $\delta=\frac{1}{20\log(T)}$. Then, for any $(\varepsilon,\delta)$-DP online learner $\mathcal{A}$ for $\mathcal{X}$ that answers $T$ queries there exists a query sequence on which $\mathcal{A}$ makes $\Omega(\log(T))$ mistakes with constant probability.
@@
Let $\mathcal{X}$ be a domain. The class of point functions over $\mathcal{X}$ admits an $(\varepsilon,\delta)$-DP online predictor with mistake bound $O(\frac{1}{\varepsilon^{2}}\log^{2}\frac{1}{\delta})$.
The proofs of Theorem 1.11 is given in Appendix A and the proof of Lemma 1.12 is given in Appendices B and C. In Appendix B we present an algorithm that satisfies Joint Differential Privacy (Kearns et al., 2015) for a version of the threshold monitor problem that applies with $k=O(\frac{1}{\varepsilon^{2}}\log^{2}\frac{1}{\delta})$ and in Appendix C we apply it to construct an online prediction algorithm.
@@ 2.2 The Formal Details
For any fixed input sequence $D\in\{0,1\}^{T}$ and $t\in[T]$, an algorithm $\mathcal{A}$ for Problem 1.6 can be fully described by a probability distribution on $[T]\cup\{\textrm{non}\}$ that corresponds to the time step $t\in[T]$ where it outputs $\top$ and halts. The event non corresponds to the algorithm outputting $\bot$ on all time steps. We use the notation $p^{D}_{t}$ for the probability that $\mathcal{A}$ outputs $\top$ (and thus halts) after the $t$-th update:
@@
1 Initialize $\ell\leftarrow 1$, $r\leftarrow T$, $D^{(0)}\leftarrow(0,0,\dots,0)$, and $c_{\mathrm{total}}\leftarrow 0$
2 for $i=1,\dots,k$ do
       $m\leftarrow\frac{1}{2}(\ell+r+1)$
        // midpoint $m$ is always an integer
3       $c_{\textrm{left}}\leftarrow\sum_{j=\ell}^{m-1}p^{D^{(i-1)}}_{j}$
4       $c_{\textrm{right}}\leftarrow\sum_{j=m}^{r}p^{D^{(i-1)}}_{j}$
5       if $c_{\mathrm{left}}\leq c_{\mathrm{right}}$ then
6             Set the $\ell$-th entry of $D^{(i-1)}$ to $1$ to obtain $D^{(i)}$
7             $c_{\mathrm{total}}\leftarrow c_{\mathrm{total}}+p_{\ell}^{D^{(i)}}$
8             $\ell\leftarrow\ell+1$
9             $r\leftarrow m-1$
10
@@ 9             $r\leftarrow m-1$
13             Set the $m$-th entry of $D^{(i-1)}$ to $1$ to obtain $D^{(i)}$
14             $c_{\mathrm{total}}\leftarrow c_{\mathrm{total}}+c_{\mathrm{left}}+p^{D^{(i)}}_{p}$
15             $\ell\leftarrow m+1$
16
@@ 3.2.2 The Key Lemma
  • First, We identify a subset $P^{(3)}$ of $P^{(2)}$ such that
    $\Pr[\mathcal{A}(S\cup\{x_{1},x_{2}\})\text{ outputs $\top$ for the $2$nd query}]$
    is almost the same (i.e., up to an error of $\eta$) for every $S\subseteq P^{(2)}$.
@@ Appendix A Lower Bound for Private Online Learning
of Theorem 1.11   Suppose for the sake of contradiction that such a learner exists and denote it by $\mathcal{A}$.
Denote $k=\sqrt{T}$, and consider the following distribution over query sequences.
  • For the $i$-th query when $i\in[1,k]$, send $(x_{i},0)$ where $x_{i}\sim\textsf{Uniform}[1,k]$ (where $\textsf{Uniform}[a,b]$ is the uniform distribution on $\{a,\ldots,b\}$).
@@
that is, the average probability, over steps $i$ in the phase, that the released hypothesis is $1$ on $k^{\prime}$ is at least $0.001/k$.
@@
Appendix B JDP Algorithm for the Mirror Problem
Input: Privacy parameters $\varepsilon\in(0,2],\delta\in(0,0.1)$, Delay parameter $K\geq 1$
@@ Appendix B JDP Algorithm for the Mirror Problem
$\varepsilon^{\prime}\leftarrow O(\varepsilon^{2}/\log(1/\delta))$; $\delta^{\prime}\leftarrow O(\delta)$
  // As in Lemma B.7
$C\leftarrow K$
for $t=1,\ldots$ do // Main Loop
       Receive an input bit $\Delta_{t}\in\{0,1\}$
       if $\Delta_{t}=1$ then // Process input $1$
            $C\leftarrow C+1$
             $r\leftarrow\textsf{Bern}(\pi_{\varepsilon^{\prime},\delta^{\prime}}(\max\{0,C-K\}))$
              // Where $(\pi(i))_{i}$ are as in Claim 3
             if $r=1$ then
                  Output $y_{t}\equiv\top$
            else
                  Output $y_{t}\equiv\bot$
             end if
      else
            Output $y_{t}\equiv\bot$
              // Input is $\Delta_{t}=0$
@@
In this section we present a construction for (a variant of) the threshold monitor problem that satisfies a weaker variant of differential privacy known as joint differential privacy (JDP) and has an error that is independent of $T$. We will use this construction in Section C as a building block of an online prediction algorithm for point functions. We first recall the definition of JDP.
Definition B.1 ($(\varepsilon,\delta)$-indistinguishable).
For two random variables $X_{1},X_{2}$ we write $X_{1}\approx_{\varepsilon,\delta}X_{2}$ if for $b\in\{0,1\}$
$\max_{S\subseteq\mathsf{supp}(X^{b})}\ln\left[\frac{\Pr[X^{b}\in S]-\delta}{\Pr[X^{1-b}\in S]}\right]\leq\varepsilon.$
We will interchangeably use notation for random variables and distributions.
Definition B.2 (Joint Differential Privacy Kearns et al. (2015)).
@@
Let $\mathcal{M}:X^{n}\rightarrow Y^{n}$ be a randomized algorithm that takes a dataset $S\in X^{n}$ and outputs a vector $\vec{y}\in Y^{n}$. Algorithm $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-joint differential privacy (JDP) if for every $i\in[n]$ and every two datasets $S,S^{\prime}\in X^{n}$ differing only on their $i$th point it holds that $\mathcal{M}(S)_{-i}\approx_{\varepsilon,\delta}\mathcal{M}(S^{\prime})_{-i}$. Here $\mathcal{M}(S)_{-i}$ denotes the (random) vector of length $n-1$ obtained by running $(y_{1},\dots,y_{n})\leftarrow\mathcal{M}(S)$ and returning $(y_{1},\dots,y_{i-1},y_{i+1},\dots,y_{n})$.
Problem B.3 (The Mirror Problem).
For a sequence of input bits $(\Delta_{t})_{t\in[T]}$ where $\Delta_{t}\in\{0,1\}$, the goal is to mirror the input, returning $y_{t}=\bot$ for $\Delta_{t}=0$ and returning $y_{t}=\top$ for $\Delta_{t}=1$. The constraints are that for a specified delay parameter $K\geq 1$:
  • For $\Delta_{t}=0$ we must return $y_{t}=\bot$.
  • We may return $y_{t}=\top$ only if $\sum_{i=1}^{t-1}\Delta_{t}\geq K$, that is, there were at least $K$ prior steps with input $\Delta_{t}=1$.
The goal is to minimize the number of mistakes. A mistake occurs at time $t$ if $\Delta_{t}=1$ and $\sum_{i=1}^{t}\Delta_{t}>K$, and $y_{t}=\bot$.
For privacy parameters $(\varepsilon,\delta)$ and delay parameter $K\geq 1$ we construct an $(\varepsilon,\delta)$-JDP algorithm for the Mirror problem. The construction is described in Algorithm 2. The algorithm uses the probabilities specified in the following claim:
Claim 3.
For $\varepsilon>0$ , $\delta\in(0,0.1)$ there are probabilities $\pi_{\varepsilon,\delta}(i)$ ($i\geq 0$) and $L\leq\lceil\frac{1}{\varepsilon}\ln((e^{\varepsilon}-1)/(2\delta))\rceil$ as follows:
  • $\pi_{\varepsilon,\delta}(0)=0$, $\pi_{\varepsilon,\delta}(1)=\delta$, $\pi_{\varepsilon,\delta}(2L)=1-\delta$
  • For $i>2L$, $\pi_{\varepsilon,\delta}(i)=1$
  • For $1\leq i<2L$, $\textsf{Bern}[\pi_{\varepsilon,\delta}(i)]\approx_{\varepsilon}\textsf{Bern}[\pi_{\varepsilon,\delta}(i+1)]$.
Proof B.4.
Let
$L=\arg\min j\ ,\delta\sum_{i=0}^{j-1}e^{i\varepsilon}>1/2\ .$
Let $\varepsilon^{\prime}\leq\varepsilon$ be the solution of $\delta\sum_{i=0}^{L-1}e^{i\varepsilon^{\prime}}=1/2$. Define
@@
Lemma B.5 (Utility of JDP-Mirror).
Algorithm 2 correctly solves the Mirror problem for delay parameter $K$ while incurring at most $2L=O(\varepsilon^{-2}\log^{2}(1/\delta))$ mistakes.
Proof B.6.
On time steps where $\Delta_{t}=0$ the algorithm always returns $y_{t}=\bot$, so there are no mistakes. On the first $K$ time steps with $\Delta_{t}=1$ we have $C\leq K$ and thus $r=0$ and hence $y_{t}=\bot$, which is as specified. At time steps where $C>K+2L$ and $\Delta_{t}=1$ we have $r=1$ and the algorithm always reports $y_{t}=\top$ so there are no mistakes. Mistakes are possible only when $\Delta_{t}=1$ and only when $K+1\leq C\leq K+2L$ so there are at most $2L=O((1/\varepsilon^{\prime})\log(1/\delta^{\prime}))$ mistakes.
Lemma B.7 (JDP of JDP-Mirror).
Algorithm 2 is $(\varepsilon,\delta)$-JDP.